OpenStudy (anonymous):

What is the solution to the equation log 2 x – log 2 4 = 5 ?

OpenStudy (dumbcow):

log_2 (4) = 2 log_2 x = 7 x = 2^7

OpenStudy (anonymous):

128

OpenStudy (dumbcow):

yep

OpenStudy (campbell_st):

use log laws log (x/4) = 5 take the exponential 2^(x/4) = 2^5

OpenStudy (campbell_st):

then x /4 = 32

OpenStudy (anonymous):

so is it 128 or 32 cus i got 128

OpenStudy (campbell_st):

lol... x/4 = 32 multiply both sides by 4 to find x x = 128

OpenStudy (anonymous):

oh lol i didnt even catch that.